Sol Summary – April 23rd

Crew 297 Sol Summary Report 23-04-2024

Sol: 09

Summary Title: Just Keep Searching

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

This morning, three crew members set foot outside of the Hab with one goal in mind: finding the NPS. They scurried across the rugged terrain with their advanced searching technologies and methodologies, diligent eyes, and unwavering dedication to locate a simple, yet elusive object. They searched and searched, giving it all they had. Yet, all of their efforts were futile. The NPS remained unfound and waiting, optimistically perhaps, to be found on another EVA on another day.

Fortunately, the crew was not left empty handed the entire day. Case the robot displayed enough capabilities today to let him out on another field deployment. With on-lookers both inside and outside the ScienceDome, Case was able to relatively smoothly (with some manual intervention) traverse an obstacle course laid out before it. Once the crew gained confidence in its abilities, they used Case to recreate a part of "The Martian": generating a hexadecimal polar grid to relay messages along.

All crew members, regardless of how successful their EVAs turned out, were treated to flavorful Southwest-style beans, rice, and veggies topped with Fritos, as prepared by Chef Laude.

Look Ahead Plan:

Tomorrow will be composed of two EVAs. The morning EVA will involve a geological expedition to the orange (via satellite imagery) regions. This EVA will also have additional testing with Case with the objective of capturing valuable 360 degree imagery of any significant geological sites. The second EVA team will do their best to evade the morning sun and find the unfound NPS from today, while scoping out suitable regions for NPS excavation research.

Anomalies in work:

Suit 11 (fan power substantially decreased mid-EVA): Battery was replaced but the air flow is still less than nominal. In-work.
Morning (data from 0900): calm winds (0 mph), chilly (59F), and sunny with clear skies
Noon: calm winds (3 mph), warm (76F), and sunny with clear skies
Afternoon (1800): strong winds (20 mph), hot (83F), and scattered cloud coverage
Crew Physical Status: Overall crew health is nominal.

EVA: There were two EVAs today. The first EVA was unsuccessful in finding the NPS after an arduous search effort. Precise coordinates have been given of the NPS and a consequent recovery EVA is planned tomorrow. The second EVA scrutinized the capabilities of Case further, and while there were setbacks, Case has been deemed ready for field deployment.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (2 EVAs)
EVA request (2 EVAs)
Support Requested: We request that the 1-hour power blackout period tomorrow be from 1600-1700. We also request support to pick-up (Qty. 2) bags of No-Burn Waste. Thanks in advance for your help.

Copyright © The Mars Society. All rights reserved. | Main Site